OpenStudy (anonymous):

7√3 + 16√3 = ? 23√3 23√6 √26 √29

OpenStudy (rational):

7+16 = ?

OpenStudy (anonymous):

23..i think i get how to do it now, its the second option right?

OpenStudy (rational):

thats a very good try but the correct answer is first option!

OpenStudy (rational):

\(\color{green}{7+16=23}\) so \( \color{green}{7}\sqrt{3} + \color{green}{16}\sqrt{3} = \color{green}{23}\sqrt{3} \)

OpenStudy (rational):

this is like adding 7 apples and 16 apples, you get 23 apples

OpenStudy (anonymous):

oooh okay thank you lots!!

OpenStudy (rational):

yw:)